OpenStudy (anonymous):

The dress store is having a sale where all merchandise is 1/4 off. A woman buys $24 of merchandise at the sale price. What would she have had to pay at the regular price?

OpenStudy (anonymous):

multiply 24 by .25 then add the number you got to 24

OpenStudy (campbell_st):

well you are told \[\frac{3}{4} = 24\] so to find the full price, divide both sides of the equation by 3/4 remember when you divide by a fraction, you have to flip the fraction and multiply so the original price is \[24 \times \frac{4}{3} = \]